tiktokstar:

(05.07) The net of a pyramid is shown below: The net of a square based pyramid, with bases labeled 7 inches and the height of the triangle labeled 14 inches. The surface area of the solid is ____ square inches.

tiktokstar:

1 attachment
tiktokstar:

help plzz

tiktokstar:

plzzzz

tiktokstar:

thanks so much

AZ:

Surface area is finding the area of each face. It's analogous to gift wrapping a box

AZ:

Her answer is wrong, by the way.

tiktokstar:

oh can you help me then

AZ:

So to find the surface area, we just have to find the area of the square and the triangles and add them all up. What is the formula for the area of a square? And the area of a triangle?

tiktokstar:

idek im so dum

AZ:

have you ever seen this formula area of a square = side * side

tiktokstar:

no

tiktokstar:

:(

AZ:

so the area of a square is side times side what is the length of the side in the image they gave you?

tiktokstar:

1 attachment
tiktokstar:

7 I think

AZ:

Yes... you already shared the image earlier. What is the length of one side of the square? Do you know what a square is?

AZ:

yes, that's good

AZ:

So area = side * side you said that the side is 7 area = 7 * 7 what is 7*7 = ?

AZ:

Just multiply the numbers @tiktokstar

tiktokstar:

ik 1 sec

tiktokstar:

49?

AZ:

Good. Now the area of a triangle = 1/2 * base * height  What is the base and what is the height?

tiktokstar:

can you just tell me the answer

AZ:

No, I cannot

AZ:

It's against the rules of the website

tiktokstar:

oh

tiktokstar:

but she did it why cant you?

AZ:

Did she give you the correct answer? No. If I give you an answer, I could be given a warning or a suspension. The purpose of this website is to help teach. If you're not able to solve this question, how will you solve the next question that's exactly like it? How will you solve the questions in the next chapter that depend on you knowing how to solve questions like this?

tiktokstar:

oh ok sorry

AZ:

Yeah, I know we all want the answer so we can move on. But if we don't learn how to do it on our own, we won't make it very far. Look at the triangle I drew and look at the triangle in your image. What is the length of the base of the triangle? Hint: It's the same length as the side of the square because the base of the triangle is right next to it

tiktokstar:

is it 14?

AZ:

That would be the height. What's the other number we have? That's our base. We need both numbers to plug into the formula for area